# Generative Modeling with Explicit Memory

### v1

@@ 5.2 Main Results
0.2 640,000 5 epochs SDE 5.81
0.2 128,000 5 epochs SDE 9.72
0.2 640,000 1 epoch SDE 5.85
0.2 1,280,000 (Full) 0 epoch SDE 4.21
0.1 640,000 5 epochs SDE 5.93
0.3 640,000 5 epochs SDE 5.68
0.2 640,000 5 epochs ODE 6.56
GMem surpasses SoTA methods in both training efficiency and effectiveness.
  1. (a)
    Pixel-space generation. For pixel-space generation, as shown in Table 1, we benchmark GMem against prior methods using FID and NI. GMem outperforms all diffusion transformer baselines and achieves competitive performance with traditional generative models, obtaining FID =1.22 on CIFAR-10 and FID =2.10 on ImageNet $64\times 64$.
  2. (b)
    Latent-space generation. For latent-space generation, in Table 1, we compare the performance and the training efficiency of GMem with other baselines on ImageNet $256\times 256$. We observe that GMem achieves a FID of 4.98 with merely $100$M NI, outperforming the current SoTA methods, including SiT and REPA. Moreover, we demonstrate that this result can be further optimized to 3.56 when extends the NI to $512$M.
GMem is a 46.7×46.7$\times$46.7 × training accelerator.
We compare the advantages of GMem in terms of training efficiency. Table 1 and Table 4 also presents the FID scores achieved by GMem with limited NI.
For the pixel-space generation, on CIFAR-10, GMem achieves performance similar to REPA with only $2.6$M NI, offering more than 3.85×3.85$\times$3.85 × speedup over REPA and over 10×10$\times$10 × speedup compared to traditional SiT.
In latent-space generation, GMem attains FID of 5.75 with NI$=64$M, surpassing the most efficient diffusion transformer baseline, REPA, which requires more than NI$=1000$M. This indicates a training speedup of over 16×. Furthermore, GMem achieves an FID of 7.26 at NI$=38.4$M, providing a speedup exceeding 46.7× compared to SiT, which achieves the similar FID $=8.3$ with NI$=1800$M.
GMem is a 10×10$\times$10 × sampling accelerator.
Table 3: FID with respect to NFE on ImageNet 256$\times$256. In this table, we present the impact of variations in NFE on FID. All models are based on SiT-L/2, trained for 100M NI (400K iterations with a batch size of 256) without using classifier-free guidance. GMem achieves higher image quality within $\texttt{NFE}=25$, demonstrating over a 10×10$\times$10 × improvement in sampling efficiency.
@@
SiT-L/2 25.6M 250 18.8
+ REPA 25.6M 250 08.4
+ GMem 25.6M 250 5.8
+ GMem 25.6M 50 7.5
+ GMem 25.6M 25 12.3
As a benefit of substantially lower FID score, GMem is capable of generating competitive samples with significantly fewer NFEs, resulting in over a 5×5$\times$5 × in sampling efficiency. Table 5.2 presents the FID scores of GMem at various NFEs.
Specifically, reducing the NFE from $250$ to $50$ leads to increase in FID of $1.7$, while still outperforming the previous SoTA, REPA, thereby achieving a sampling speedup of more than 5×5$\times$5 ×. Additionally, decreasing the NFE from $250$ to $25$ allows GMem to generate samples that remain competitive with those of SiT, attains an FID of 12.3, providing more acceleration of more than 10×10$\times$10 ×.
GMem is NOT a memory machine.
Storing information related to source data often raises privacy and copyright concerns, particularly in the context of image generation [58, 64, 5]. Although our approach incorporates a memory bank $\mathbf{M}$, we clarify that GMem does not function as a system solely for storing pixel-level detailed information from images. Instead, the memory bank stores high-level semantic representations, enabling $\mathbf{M}$ to serve as an external knowledge repository. This design facilitates faster training and sampling processes while mitigating privacy considerations.
Figure 4: Diverse image generation. The figure presents selected samples from ImageNet $256\times 256$ generated by the SiT-XL/2 + GMem model using checkpoints with $\texttt{NI}=512M$ and without classifier-free guidance. The top row shows the original image corresponding to the memory snippet, followed by four rows of images generated through different noises when sampling. For instance, in the sixth column, GMem adeptly captures the semantic essence of a castle nightscape from memory snippets, while presenting variations in elements such as appearance, lighting, perspective, and other nuanced details.
We demonstrate the semantic information within the memory bank $\mathbf{M}$ through experiments conducted on ImageNet $256\times 256$. Specifically, for one memory snippet $\mathbf{s}$ in $\mathbf{M}$, we apply different noise during the sampling process and compare with the generated samples. As shown in Figure 4, GMem generates images that are semantically aligned with the original input while differ in specific details. This outcome confirm that GMem does not merely memorize $\mathbf{M}$ or learn a direct mapping from $\mathbf{M}$ to $\mathcal{Z}$, but instead effectively leverages the high-level semantic knowledge stored in $\mathbf{M}$ during the image generation process.
Appendix D further demonstrates GMem’s flexibility in capturing the knowledge from $\mathbf{M}$, where we interpolate between representations from two different classes to generate a series of smoothly transitioning images.
@@ 5.3 Ablation Study
The effectiveness of GMem depends on several key factors, including dropout ratio, bank size, bank training settings, solver, and backbone architectures. In this subsection, we set NI to $100M$ and conducted a series of ablation studies on ImageNet $256\times 256$. We systematically examined these factors and concluded that, although each variable has a slight impact on the optimal performance, GMem still consistently generates high-quality images efficiently. The results are reported in Table 5.2 and Table 4.
Dropout ratio of memory snippets.
We first investigate the impact of different dropout ratios on model performance. As shown in Table 5.2, introducing a moderate level of dropout proves advantageous, with a 20% rate applied to the memory snippets yielding the best results. We argue that this dropout mechanism implicitly reduces the model’s reliance on memory snippets, thus improving generalizability and overall performance.
Configurations of building memory bank.
We examine the impact of various memory bank configurations on performance, we reported the results in Table 5.2. Our findings indicate that larger memory banks generally yield superior results, with bank size $=1{,}280{,}000$ leading to optimal result of FID $=4.21$. We propose that excessively small memory banks impede the retention of essential knowledge, thereby degrading FID performance, while overly large memory banks increase storage requirements and computational costs. Typically, allocating 50% of total images in the dataset to the memory bank—approximately 1.8 GB for ImageNet $256\times 256$—provides satisfactory outcomes. Additionally, we demonstrate that training the memory bank for five epochs is sufficient to achieve the optimal results.
Choice of SDE/ODE solver.
We observe that SDE solver consistently produces higher quality images than the ODE solver, achieving a FID score reduction of 0.25, as shown in  Table 5.2. Consequently, we employ SDE solvers in all main experiments.
Figure 5: Ablation study on the variation of FID-50K with respect to NI. We analyze the impact of NI during the training of GMem using different backbones on CIFAR-10 dataset. Specifically, while GMem-XL initially shows a higher FID than GMem-L at the early stages of training, after $1.75$M NI, GMem-XL consistently outperforms both GMem-L and GMem-B in terms of FID and maintains this lower value until convergence, where the final FID reaches $1.22$.
Neural network architecture.
Table 4: FID comparisons with various SiT backbones. In this table, we report the FID of GMem with different backbones on ImageNet $256\times 256$ at various NI values. For comparison, we also present the performance of the state-of-the-art baseline REPA at similar NI or comparable FID values. Note that $\downarrow$ indicates that lower values are preferred and all results reported are without classifier-free guidance.
[…]
SiT-B/2 130M 100M 33.0
+ REPA 130M 100M 24.4
+ GMem 130M 100M 15.4
SiT-L/2 458M 100M 18.8
+ REPA 458M 180M 8.4
+ GMem 458M 100M 5.8
SiT-XL/2 675M 100M 17.2
+ REPA 675M 100M 7.9
+ GMem 675M 100M 5.0
SiT-XL/2 675M 1800M 8.3
+ GMem 675M 38.4M 7.3
+ REPA 675M 1024M 5.9
+ GMem 675M 64M 5.8
+ GMem 675M 512M 3.6
We delve into the scalability of the neural network architecture of GMem by varying its size. Table 4 presents the FID-50K scores of GMem across different model sizes on ImageNet $256\times 256$. Our findings reveal that larger models not only converge more rapidly but also achieve lower FID scores, aligning with the observations of diffusion transformers by Yu et al. [70] and Ma et al. [44]. Notably, compared to smaller models, GMem-XL consistently reaches the same FID levels more quickly and ultimately converges to a lower FID of 3.56 on ImageNet $256\times 256$.
This scalability trend extends to pixel-space generation. For example, on CIFAR-10, as depicted in  Figure 5, GMem-XL achieves an FID of 1.22 more efficiently than smaller model variants. These results underscore the scalability of GMem, demonstrating its ability to maintain high generation quality while enhancing training and sampling efficiency across diverse architectures and datasets.
5.4 Additional Applications
Transferable memory bank across datasets.
Appendix E demonstrates the transferability and generalization of the memory bank used to guide GMem across different datasets. Specifically, we train $\textsc{GMem}_{\textrm{IN64}}$ and $\textsc{GMem}_{\textrm{CIFAR}}$ models on ImageNet $64\times 64$ and CIFAR-10, respectively, resulting in memory banks of $\mathbf{M}_{\textrm{IN64}}$ and $\mathbf{M}_{\textrm{CIFAR}}$. We then directly apply $\mathbf{M}_{\textrm{IN64}}$ to guide the sampling process of $\textsc{GMem}_{\textrm{CIFAR}}$. Our results show that $\textsc{GMem}_{\textrm{CIFAR}}$ is still able to generate information consistent with the knowledge provided by $\mathbf{M}_{\textrm{IN64}}$.
Application to T2I generation.
This integration is achieved through a two-step process:
  1. (a)
    Pretraining the mapping function: We pretrain a mapping function $\boldsymbol{\psi}$ that transforms text prompt distributions $\mathbf{p}({\text{text}})$ into memory snippet distributions $\mathbf{p}(\mathbf{s})$ using a simple contrastive loss [54].
  2. (b)
Comprehensive descriptions of the T2I generation process, experimental setup, and the resulting generated images are provided in Appendix E.
@@ 6 Conclusion
In this paper, we introduce a new paradigm for diffusion models, GMem, which accelerates both training and sampling processes by incorporating a memory bank. Our method achieves SoTA performance on CIFAR-10 and ImageNet $64\times 64$, improving training efficiency by more than 10$\times$ and sampling efficiency by over 5$\times$ across datasets. Furthermore, we demonstrate that our memory bank can be transferred across different datasets and can be easily adapted to T2I tasks by generating a memory bank, further showcasing the flexibility and generalization capability of GMem.
Supplementary Material
We include additional analysis, experiments details, and results of downstream tasks in this supplementary material. In Section A, we provide a detailed analysis and experimental settings of the memory bank. In Section B, we present additional experiment, including the ablation of different encoders. In Section C, we offer further implementation details on the training settings of GMem on the CIFAR-10 dataset, ImageNet 64$\times$64 dataset, and ImageNet 256$\times$256 dataset. In Section D, we provide detailed analysis on the interpolation experiments. In Section E, we demonstrate that the bank learned from one dataset can be easily transferred to another dataset without any additional training. Besides, in Section F, we also show that GMem can be seamlessly applied to text-to-image generation tasks. Finally, in Section G, we supplement the derivation of the objective functions for flow-based and diffusion-based models.
Appendix A Memory Bank
In this section we provide a detailed analysis of the memory bank, including the theoratic analysis, building details of the memory bank and the ablative study of the memory bank size related to performance on downstream tasks.
A.1 Analysis
To ensure that for every input $\mathbf{x}$ sampled from the data distribution $D$, the normalized feature $\nicefrac{{\boldsymbol{f}(\mathbf{x})}}{{\|\boldsymbol{f}(\mathbf{x})\|}}$ corresponds exactly to a vector $\mathbf{s}$ present in the matrix $\mathbf{M}$, we require that:
$\underset{\mathbf{s}\in\mathbf{M}}{\min}\left\{\|\mathbf{s}-\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\|\right\}=0,\,\forall\mathbf{x}\sim D\,.$ (11)
Since we enforce $\|\mathbf{s}\|=1$, the condition in (11) implies that:
$\underset{\mathbf{s}\in\mathbf{M}}{\max}\left\{\|\mathbf{s}\cdot\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\|\right\}=1,\,\forall\mathbf{x}\sim D\,.$ (12)
Consequently, our objective is to satisfy:
${\max}\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}=1,\,\forall\mathbf{x}\sim D\,.$ (13)
This leads us to formulate the optimization problem as follows:
$\operatorname*{arg\,min\,}_{\mathbf{M}\in\mathbb{R}^{n\times m}}\left\{\left(\mathbb{E}_{\mathbf{x}\sim D}\left[\max\left\{\mathbf{M}\frac{\boldsymbol{f}(\mathbf{x})}{\|\boldsymbol{f}(\mathbf{x})\|}\right\}-1\right]\right)^{2}\right\}\,.$ (14)
A.2 Setting
As described in Table 6 and Table 7, across all experiments, we set bank unit $m$ to 768 to match the size of the output hidden dimension of the encoder. Additional, we set the memory bank size $n$ to 50k for CIFAR-10 and 1.2M for ImageNet 64$\times$64 and ImageNet 256$\times$256.
When constructing the memory bank, we use the Euclidean distance as the similarity measure between the features output by encoder and the memory bank and use MSE as loss. We choose the Adam optimizer with a learning rate of 1e-2, $\beta_{1}$ of 0.9, and $\beta_{2}$ of 0.999. The memory bank is trained for 80 epochs on CIFAR-10 with the batch size of 64.
A.3 Building
We first detailed the building process of the memory bank. Initially, the memory bank is populated with features and then fine-tuned using the outputs from the encoder. Specifically, during fine-tuning, we compute batched features from both the encoder and the memory bank. We then calculate the similarity between these batched features and the snippets within the memory bank, and compute the loss based on the most similar snippet. Finally, we update the memory bank features according to the computed similarities.
Table 5: Bank ablation study. We also conduct an ablation study with respect to the bank size on CIFAR-10. Note that $\downarrow$ indicates that lower values are preferred.
[…]
SiT-XL/2
+ (GMem) 50K 22.4M 1.22
+ (GMem) 25K 22.4M 1.50
+ (GMem) 5K 22.4M 3.90
We present the results of our ablation study on memory bank size in  Table 5. Overall, a larger memory bank leads to better performance, as indicated by lower FID scores. Specifically, when the bank size is 50,000—equivalent to the entire dataset—the performance is optimal. Reducing the bank size to 25k results in a slight performance decrease to FID $=1.50$, yet it still surpasses the baseline. However, further reducing the bank size to 5k significantly degrades performance, indicating the importance of maintaining a relatively large memory bank for optimal results.
Overall, GMem requires less than 150MB of additional storage (the size of a 50K memory bank) to achieve a 10$\times$ faster training speed and a 5$\times$ faster inference speed on downstream tasks.
Appendix B Supplementary Experiments
@@
B.1 Ablation Study on the Encoder
We conduct an ablation study to investigate the impact of different encoders on the performance of GMem. Specifically, we choose the visual encoder from CLIP [54] as an alternative to show that GMem can be applied across encoders.
@@ Results
Appendix C Implementation Details
C.1 Diffusion transformer architecture
We closely follow the architecture used in REPA [70] and SiT [44]. Similar to a Vision Transformer [13], In this architecture, the input image is divided into patches, reshaped into a one-dimensional sequence of length $N$, and then processed by the model. Unlike the original SiT, REPA includes additional modulation layers called AdaIN-zero layers at each attention block. These layers scale and shift each hidden state based on the given timestep and additional conditions.
For latent space generation, similar to REPA, our architecture uses a downsampled latent image $z=E(x)$ as input, where $x$ is an RGB image and $E$ is the encoder of the Stable Diffusion Variational Autoencoder (VAE) [55]. For pixel space generation, we remove the encoder and directly use the RGB image as input. Specifically, we modify the original SiT by changing the number of channels from 4 to 3 and directly feed the transformed RGB image into the model. For ImageNet $64\times 64$, as an example of pixel space generation, we aslo adjust the patch size of SiT from 4 to 2 to maintain a consistent sequence length of $N=256$ patches.

### v2

@@ 5.2 Training and sampling efficiency of GMem
DiT-XL 675M 1400 9.62
SiT-XL 675M 1400 8.61
  + REPA 675M 800 5.90
LightningDiT-XL 675M 800 2.17
  + GMem 675M 28 7.66
  + GMem 675M 32 4.86
  + GMem 675M 160 1.53
GMem is a $\mathbf{50\times}$ training accelerator.
We evaluate training efficiency and performance across pixel-space and latent-space generation tasks. As shown in Table 1 and Table 5.2, GMem achieves competitive FID scores with significantly fewer training epochs.
Pixel-space generation: On CIFAR-10 in Table 1, GMem matches REPA’s performance in just $450$ epochs, delivering a 3.85×3.85$\times$3.85 × speedup over REPA and over 10×10$\times$10 × compared to traditional SiT.
Latent-space generation: On ImageNet in Table 1, GMem achieves FID $=1.53$ on ImageNet $256\times 256$ and $1.71$ on ImageNet $512\times 512$ in just $160$ and $400$ epochs, without using classifier-free guidance. As shown in Table 5.2, GMem achieves competitive FID scores with significantly fewer epochs than the most efficient baselines REPA and SiT: (1) With only $32$ epochs ($\mathbf{25\times}$ speedup), GMem achieves FID $=4.86$, outperforming REPA, which requires $800+$ epochs. (2) With only $28$ epochs ($\mathbf{50\times}$ speedup), GMem reaches FID $=7.66$, surpassing SiT’s FID $=8.61$ at $1{,}400$ epochs.
GMem is a $\mathbf{10\times}$ sampling accelerator.
Table 3: GMem accelerates sampling by 10$\times$. This table examines the impact of different ${NFE}$ choices on ${FID}$. For a fair comparison with REPA, we use SiT-L/2 as backbone and trained GMem for ${Epochs}=20$ (400K iterations with a batch size of 256) without using classifier-free guidance. $\downarrow$ means lower is better.
@@
SiT-L/2 20 250 18.8
  + REPA 20 250 08.4
  + GMem 20 250 5.8
  + GMem 20 50 7.5
  + GMem 20 25 12.3
By generating competitive samples with significantly fewer NFE, GMem improves sampling efficiency by over 5×5$\times$5 ×. Table 5.2 shows the FID scores of GMem across varying NFEs and baselines: (1) With NFE $=50$ ($\mathbf{5\times}$ speedup), GMem outperforms REPA with NFE $=250$. (2) With NFE only $25$ ($\mathbf{10\times}$ speedup), GMem still surpasses SiT with NFE $=250$.
5.3 Manipulating Memory for Diverse Images
Diverse image generation.
We first demonstrate in Figure Figure 5 that even for memory snippets present in the training set, GMem can generate images that deviate significantly from the training data. This illustrates that GMem does not merely memorize images from the training set, but instead leverages the network’s learned generalization capabilities to produce novel samples. Specifically, when a memory snippet corresponding to the semantic concept of a "lighthouse" is provided, the generated image of the lighthouse exhibits clear differences in background, color, and other attributes compared to the most similar image in the training set. Similarly, when a memory snippet representing an "orange" is given, GMem generates an image of an orange placed on a plate, showing distinct differences in presentation from the close-up images of oranges in the training set.
Novel image generation.
The decoupled memory bank allows GMem to generate entirely new classes or samples without additional training. By integrating snippets via SVD-based methods (see Section 4.3), GMem can embed external knowledge into its memory bank and synthesize previously unseen categories. For instance, as shown in Figure 4, providing a “low-poly” style snippet lets GMem produce images exhibit “low-poly” characteristics but distinct from the original input—even if such a style was never in the training data. This training-free style transfer highlights how GMem can leverage memory snippets to create novel samples and underscores the flexibility of its decoupled memory bank in adapting to new styles or categories without retraining.
Compositional image generation.
Next, we will explore generating new concepts by manipulating existing snippets through latent space interpolation. Rather than introducing a new snippet, we linearly combine two snippets stored in $\mathbf{M}$ and the interpolated snippet produces images blending features from both “parent” snippets, creating hybrid or novel concepts. For example, interpolating between a “dog” and a “hat” snippet generates a new concept “dog wearing a hat”. This approach also allows introducing artistic styles to existing classes, e.g., we can generate a “swan swimming in a green river” by interpolating between a internal snippet “swan” and a external knowledge “green background” snippet. These results highlight GMem’s ability to composite internal and external knowledge for generating new concepts.
@@ 5.4 Ablation Study
The performance of GMem depends on several factors: bank size, backbone architecture, solver, and masking strategies. We conduct ablation studies on ImageNet $256\times 256$ with epochs = $64$, and found that while each factor slightly impacts optimal performance, GMem consistently generates high-quality images efficiently.
Optimal masking ratio for memory bank.
We investigate the impact of varying masking ratio on model performance. As illustrated in Table A.2, a moderate mask ratio of $0.4$ balances performance and generalization, reducing reliance on memory snippets while avoiding performance degradation from excessive masking. This configuration achieves the best performance, yielding an FID $=5.70$.
We also evaluate different masking strategies outlined in Appendix A.2. GMem adapts effectively to all three strategies, demonstrating robustness to noisy memory bank conditions.
SVD-based decomposing strategy enhances generalization.
We investigate the impact of memory bank compression methods, with results summarized in Table A.2. Applying the decomposing strategy from Section 4.3 while halving the memory bank size reduces FID by approximately $0.15$, likely due to increased sample diversity from slight noise. Additionally, we explore snippet-level compression by randomly selecting $640$ samples per class, resulting in a half-sized memory bank. This reduction degrades FID performance by only $0.02$, highlighting substantial redundancy in the memory bank and underscoring the necessity of compression.
Additional findings.
We show that SDE solver is always better than ODE solver in Appendix A.2 and Table A.2. We also find that larger backbones show better performance and GMem generalize well with various backbones and visual tokenizers in Appendix A.3 and Table 5.
@@ 6 Conclusion
In this paper, we introduce a new paradigm for diffusion models, GMem, which accelerates both training and sampling processes by decoupling the memorization from the model backbone into a memory bank. GMem achieves SoTA performance on CIFAR-10 and ImageNet $256\times 256$ and $512\times 512$ and improving training efficiency by more than 50$\times$ and sampling efficiency by over 10$\times$ on ImageNet $256$. Additionally, the decoupled memory bank supports training-free adaptation to new images not present in the training set.
Acknowledgements
This work is supported in part by the Research Center for Industries of the Future (RCIF) at Westlake University, Westlake Education Foundation, and Westlake University Center for High-performance Computing.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Supplementary Experiments
@@
A.1 Ablation Study on Vision Encoder
We conduct an ablation study to investigate the impact of different encoders on the performance of GMem. Specifically, we choose the visual encoder from CLIP (Radford et al., 2021) as an alternative to show that GMem can be applied across encoders.
@@ Results
A.2 Ablation Study on Memory Bank Size, Masking Strategies and Solver
Table 4: Ablation study and sensitive analysis. All models use FasterDiT-B, trained for ${Epochs}\!=\!64$ w/o using classifier-free guidance on ImageNet $256\times 256$. $\downarrow$ means lower is better.
[…]
1.2B Zero 0.4 SDE 5.70
1.2B $\times$ Zero 0.4 SDE 5.85
640K Zero 0.4 SDE 5.72
1.2B Noise 0.4 SDE 6.79
1.2B Random 0.4 SDE 6.62
1.2B Zero 0 SDE 6.28
1.2B Zero 0.3 SDE 5.75
1.2B Zero 0.4 ODE 6.70
Additional Masking Strategies.
We also explored two other masking strategies: random mask and noise mask. Specifically, random mask replaces a randomly selected portion of each batch with Gaussian noise, while noise mask adds noise to the entire memory snippet. The results for these two masking strategies are presented in Table A.2. We found that zeroing out part of the snippet (the Zero mask strategy) consistently performed best across all experiments. Therefore, we adopted Zero mask for all major experiments.
SDE solver is superior.
SDE solvers consistently outperform ODE solvers, reducing FID by $1.0$ ( Table A.2). Thus, SDE solvers are used in all main experiments.
A.3 Ablation Study on Model Backbones and Tokenizers
Larger architectures excel.
Table 5: GMem consistently generates high-quality samples across different backbones and image tokenizers. This table reports the FID of GMem with different backbones and visual tokenizers on ImageNet $256\times 256$. For a fair comparison, we train all models for 64 epochs. $\downarrow$ means lower is better and all results reported are without classifier-free guidance.
[…]
SiT-B SD-VAE 130M 64 22.25
SiT-L SD-VAE 458M 64 6.49
SiT-XL SD-VAE 675M 64 6.31
LightningDiT-B VA-VAE 130M 64 5.70
LightningDiT-B DC-AE 130M 64 5.97
We examine the scalability of GMem by testing various model sizes and architectural configurations. Table 5 presents the FID-50K scores of GMem across different model sizes on ImageNet $256\times 256$: larger models not only converge faster but also achieve lower FID. This trend aligns with findings from Yu et al. (2024) and Ma et al. (2024) on diffusion transformers and extends to pixel-space generation. For example, on CIFAR-10, GMem-XL achieves an FID $=1.22$, outperforming smaller model variants efficiently, as depicted in  Figure 9.
GMem generalize well with various backbones and visual tokenizers.
We compare different backbone architectures and visual tokenizers in Figure 5. LightningDiT consistently outperforms SiT under identical configurations, corroborating findings from Yao & Wang (2025). Additionally, DC-AE and VA-VAE tokenizer yields better results than SD-VAE , likely due to their larger parameter capacity.
Appendix B Implementation Details
B.1 Diffusion transformer architecture
We closely follow the architecture used in REPA (Yu et al., 2024) and SiT (Ma et al., 2024). Similar to a Vision Transformer (Dosovitskiy et al., 2021), In this architecture, the input image is divided into patches, reshaped into a one-dimensional sequence of length $N$, and then processed by the model. Unlike the original SiT, REPA includes additional modulation layers called AdaIN-zero layers at each attention block. These layers scale and shift each hidden state based on the given timestep and additional conditions.
For latent space generation, similar to REPA, our architecture uses a downsampled latent image $z=E(x)$ as input, where $x$ is an RGB image and $E$ is the encoder of the Stable Diffusion Variational Autoencoder (VAE) (Rombach et al., 2022). For pixel space generation, we remove the encoder and directly use the RGB image as input. Specifically, we modify the original SiT by changing the number of channels from 4 to 3 and directly feed the transformed RGB image into the model. For ImageNet $64\times 64$, as an example of pixel space generation, we aslo adjust the patch size of SiT from 4 to 2 to maintain a consistent sequence length of $N=256$ patches.